### v1

# Predicting Case Suffixes With Activity Start and End Times: A Sweep-Line Based Approach

Abstract
Predictive process monitoring techniques support the operational decision-making by predicting future states of ongoing cases of a business process. A subset of these techniques predict the remaining sequence of activities of an ongoing case (case suffix prediction). Existing approaches for case suffix prediction generate sequences of activities with a single timestamp (e.g. the end timestamp). This output is insufficient for resource capacity planning, where we need to reason about the periods of time when resources will be busy performing work. This paper introduces a technique for predicting case suffixes consisting of activities with start and end timestamps. In other words, the proposed technique predicts both the waiting time and the processing time of each activity. Since the waiting time of an activity in a case depends on how busy resources are in other cases, the technique adopts a sweep-line approach, wherein the suffixes of all ongoing cases in the process are predicted in lockstep, rather than predictions being made for each case in isolation. An evaluation on real-life and synthetic datasets compares the accuracy of different instantiations of this approach, demonstrating the advantages of a multi-model approach to case suffix prediction.
1 Introduction
Predictive Process Monitoring (PPM) techniques provide runtime insights to operational managers by exploiting models trained on event logs to predict the future states of ongoing cases of a process. For example, PPM techniques can predict the outcome of ongoing cases (e.g. will a customer accept or reject a product?) [21], the remaining time of ongoing cases [22], the next activity in a case, or the sequence of remaining activities in a case (a.k.a. the case suffix[16].
Existing approaches for case suffix prediction generate sequences of activities with a single timestamp (e.g., end timestamp). This is insufficient for use cases such as capacity planning and scheduling, where managers need both start and end timestamps to calculate processing and waiting times, assess resource utilization, and determine whether capacity is adequate for the expected workload.
In this setting, we propose an approach for predicting case suffixes composed of activities with start and end timestamps. Our method employs a sweep-line-based technique [14], where suffixes for all ongoing cases are predicted collectively, rather than individually. The approach follows a three-stage prediction process. First, a model predicts the next activity. Given this activity, a second model estimates the inter-start time, i.e., the time elapsed between the start of the previous activity and the start of the predicted activity. Finally, a third model predicts the processing time i.e., the duration required to complete the activity once it has started. Using these predictions, we derive the start and end timestamps of each predicted activity instance in the case suffix. The start timestamp is obtained by adding the inter-start time to the start timestamp of the previous activity instance, while the end timestamp is derived by adding the processing time to the start timestamp of the predicted activity instance.
We report on an evaluation that compares the accuracy of our multi-model approach, along control-flow and temporal metrics, relative to approaches that predict the next activity and timestamp via a single model approach.
2 Related Work
Several studies have addressed the problem of predicting case suffixes via deep learning. Tax et al. [19] and Evermann et al. [6] proposed LSTM-based architectures to predict the next activity, the case suffix, and the remaining time until case completion, using one-hot and embedded sequence encodings, respectively. Camargo et al. [4] extended these approaches by proposing multi-output LSTM models capable of jointly predicting activity, end timestamp, and roles. Taymouri et al. [20] proposed a GAN-based approach, while Wuyts et al. [24] and Rama-Maneiro et al. [17] introduced transformer and attention-based encoder-decoder models to predict full suffixes in a single step. Ketykó et al. [10] highlight that suffix prediction performance is influenced by both architecture and sampling/decoding strategies. Rama-Maneiro et al. [15] propose DOGE, a DRL-based sampler that replaces traditional methods like Argmax with learned policies. Pasquadibisceglie et al. [12] proposed CNN-based case suffix predictors, and later extended this approach using Large Language Models (LLMs) [13].
The above approaches follow a single model (SM) paradigm, which jointly predicts the next activity and one timestamp (either the start or the end time). None of these approaches is designed to estimate both the start and the end timestamps of the activities in the case suffix. In other words, none of these approaches sepatrates the waiting time and the processing time of the activities in the suffix. Our approach tackles this limitation by departing from the SM approach and, instead, training three separate models individually to predict the next activity, its inter-start time, and its processing time. This approach allows us to flexibly calculate waiting times and processing times separately.
Other researchers have combined simulation with learning-based techniques. Camargo et al. [3] and Meneghello et al. [11] proposed integrating LSTM predictors with discrete-event simulation models derived from discovered process models. These methods simulate synthetic traces, including timestamps, under hypothetical “what-if” scenarios and can be used to evaluate process changes or capacity modifications. In contrast, our approach focuses on generating case suffixes for “as-is” processes by leveraging real-time event logs without assuming process changes or relying on synthetic generation.
3 Approach
The proposed approach consists of two phases: an offline phase for training predictive models for case suffix prediction and an online phase which adopts a sweep-line based method wherein the suffixes of all ongoing cases in the process are predicted in lockstep, rather than predictions being made in isolation.
3.1 Offline Phase
The offline phase (Fig.1) focuses on preparing the training data and learning predictive models necessary for case suffix prediction. This phase begins with transforming an event log into structured input sequences by extracting both intra-case features (capturing case-specific temporal patterns), and inter-case features (capturing system-wide dynamics, such as resource utilization and workload). These features are encoded and assembled into fixed-size sequences using an n-gram strategy to standardize inputs for model training. We train three specialized BiLSTM-based models – one each for predicting the next activity($\alpha$), inter-start time($\beta$), and processing time($\gamma$). Below, we describe the offline phase.
Figure 1: Offline Phase
3.1.1 Input:
To train models for case suffix prediction, we take as input a type of event log called an activity instance log. Table 1 contains an excerpt of such a log. Each row corresponds to an activity instance. For each activity instance, the proposed approach requires a case ID, activity name, start and end timestamps. Every row must have a value for each of these four attributes. However, for some activity instances (e.g., the second one in Case 2) the value of the end timestamp may be null (denoted as $\emptyset$). When an activity instance has a null end timestamp, it means that it has not yet completed. Our approach does not require other attributes in the log. In particular, it does not require a resource attribute nor a role attribute. On the other hand, the approach does assume that the case ID, activity name, start and end timestamps are accurate, otherwise the machine learning models would be trained to produce inaccurate case suffixes. Formally, we define an activity instance log as follows.
Definition 1(Activity Instance Log)
An Activity Instance Log $\mathcal{L}$, is a finite set of records $r\in\mathcal{L}$, where each record $r=(c,a,T^{\text{start}},T^{\text{end}})$ consists of the following elements: $c\in C$,where $C$ is the set of unique case identifiers for process instances; $a\in A$, where $A$ is the set of possible activities; and $T^{\text{start}},T^{\text{end}}\in\mathbb{T}\cup\{\emptyset\}$, representing the start timestamp and end timestamp of an activity, with $\mathbb{T}$ being the domain of time values and $\emptyset$ denotes a missing (null) timestamp.
Table 1: Activity Instance Log
Case ID Activity $T^{\text{start}}$ $T^{\text{end}}$
1 Received Query 08:00 08:05
Assigned to Rep 08:10 08:15
Query Resolved 08:20 09:00
Cust Notified 09:05 09:10
2 Received Query 09:15 09:20
Assigned to Rep 09:25 $\emptyset$
3 Received Query 10:00 10:05
Assigned to Rep 10:10 10:15
Query Resolved 10:20 11:00
The Activity Instance Log ($\mathcal{L}$) may be structured as a set of activity instance traces ($\sigma_{c}$), where each trace is an ordered sequence of activities for a specific case identifier ($c$). An activity instance trace is defined as follows.
Definition 2(Activity Instance Trace of Log $\mathcal{L}$)
Given an activity instance log $\mathcal{L}$, an activity instance trace $\sigma_{c}$ of $\mathcal{L}$ is a sequence of activity instances $\sigma_{c}=\langle r_{c,1},r_{c,2},\dots,r_{c,n}\rangle$, such that:
  • All activity instances $r_{c,i}$ in $\sigma_{c}$ share the same case identifier $c$.
  • Every activity instance in $\mathcal{L}{}$ that has $c$ as its case identifier is part of $\sigma_{c}$.
  • The activity instances in $\sigma_{c}$ are chronologically ordered by start timestamp.
Herein, we write $\sigma_{c}\in\mathcal{L}{}$ to denote that $\sigma_{c}$ is a trace of log $\mathcal{L}{}$.
Our approach takes as input an activity instance log $\mathcal{L}$ generated by a collection of cases over a specified time period and a time point called cutoff denoted by ($t_{\text{cutoff}}$). Given $\mathcal{L}{}$ and $t_{\text{cutoff}}$, we derive the training log, denoted as $\mathcal{L}_{train}$. The training log includes all $\sigma_{c}{}\in\mathcal{L}{}$ that fulfill the following two conditions:
  • Every activity instance in $\sigma_{c}{}$ has a non-null end timestamp, i.e., $\neq$ $\emptyset$.
  • Every activity instance in $\sigma_{c}{}$ has an end timestamp $\leq$ $t_{\text{cutoff}}$.
To train a deep learning model, we enhance each case in the training log with intra- and inter-case features. Alg. 1 describes the procedure for enhancing a given trace $\sigma_{c}{}$ with such features. This procedure is applied to each $\sigma_{c}{}$ $\in$ $\mathcal{L}_{train}$.
Input: An activity instance log $L$, An activity instance trace $\sigma_{c}\in L$
Output: An enhanced activity instance $\sigma_{c}$ including inter- and intra-case features
1
2foreach trace prefix $P=\langle r_{c,1},\dots,r_{c,k}\rangle$ in $\sigma_{c}$ do
3  for $i\leftarrow 1$ to $k$ do
4     Let $r_{c,i}=(c,a_{c,i},T^{\text{start}}_{c,i},T^{\text{end}}_{c,i})$;
5     $\delta^{\text{proc}}_{c,i}\leftarrow T^{\text{end}}_{c,i}-T^{\text{start}}_{c,i}$;
6     if $i=1$ then
7        $\delta^{\text{start}}_{c,1}\leftarrow 0$;
8
9    else
10        $\delta^{\text{start}}_{c,i}\leftarrow T^{\text{start}}_{c,i}-T^{\text{start}}_{c,i-1}$;
11
12    $r_{c,i}\leftarrow r_{c,i}\oplus(\delta^{\text{proc}}_{c,i},\delta^{\text{start}}_{c,i})$
13
14 Let $T^{\text{eval}}\leftarrow T^{\text{start}}_{c,k}$;
15
16 BusyResources $\leftarrow\emptyset$;
17  TotalActiveInstances $\leftarrow 0$;
18
19 foreach $r=(c^{\prime},a^{\prime},T^{\prime}_{\text{start}},T^{\prime}_{\text{end}})\in L$ do
20     if $T^{\prime}_{\text{start}}\leq T^{\text{eval}}\leq T^{\prime}_{\text{end}}$ then
21        Increment TotalActiveInstances;
22        if $a^{\prime}=a_{c,k}$ then
23           Add $r$ to BusyResources;
24
25
26
27 $\text{WIP}_{c,k}\leftarrow$ TotalActiveInstances;
28  $r_{c,k}\leftarrow r_{c,k}\oplus(\text{WIP}_{c,k})$;
29  $\text{Utilization}_{c,k}\leftarrow|\texttt{BusyResources}|$;
30  $r_{c,k}\leftarrow r_{c,k}\oplus(\text{Utilization}_{c,k})$;
31  Let $\tau\leftarrow 0.2\times\text{log duration}$;
32  Initialize RecentCaseStarts $\leftarrow\emptyset$;
33  foreach case $c^{\prime}\in L$ do
34     Let $T^{\text{start}}_{c^{\prime}}$ be the first start timestamp in $\sigma_{c^{\prime}}$;
35     if $T^{\text{start}}_{c^{\prime}}\geq T^{\text{eval}}-\tau$ then
36        Add $c^{\prime}$ to RecentCaseStarts;
37
38
39 $\lambda_{c,k}\leftarrow|\texttt{RecentCaseStarts}|/\tau$;
40  $r_{c,k}\leftarrow r_{c,k}\oplus(\lambda_{c,k})$;
41
Algorithm 1 Enhancing an activity instance trace $\sigma_{c}$ with Inter & Intra-case Features
3.1.2 Intra-Case Features:
We extract features from each trace prefix of a trace $\sigma_{c}$. Given a trace $\sigma_{c}=\langle r_{c,1},\dots,r_{c,n}\rangle$ capturing the execution of case $c$, and given an index $1\leq k\leq n$, the trace prefix of case $c$ at index $k$, denoted $\text{$\mathcal{P}$}(c,k)$, is the sequence $\langle r_{c,1},\dots,r_{c,k}\rangle$. For example, in Table 1, the prefix of case 1 at index 2 $\text{$\mathcal{P}$}(1,2)=\langle(\text{Received Query},08{:}00,08{:}05),(\text{Assigned to Rep},08{:}10,08{:}15)\rangle$.
For each activity instance $r_{c,k}$ = $(c,a_{c,k},T^{\text{start}}_{c,k},T^{\text{end}}_{c,k})$ in a prefix, we construct an enhanced activity instance $(c,a_{c,k},T^{\text{start}}_{c,k},T^{\text{end}}_{c,k},\delta^{\text{start}}_{c,k},\delta^{\text{proc}}_{c,k})$, where $\delta^{\text{start}}_{c,k}=T^{\text{start}}_{c,k}-T^{\text{start}}_{c,k-1}$ denotes the inter-start time, and $\delta^{\text{proc}}_{c,k}=T^{\text{end}}_{c,k}-T^{\text{start}}_{c,k}$ is the processing time. For the first activity, we define $\delta^{\text{start}}_{c,1}=0$. The procedure for extracting vector $v$ is implemented in lines (1-8) of Alg. 1. The algorithm iterates over each trace $\sigma_{c}\in\mathcal{L}$, where $\mathcal{L}$ is the activity instance log grouped by case. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, it processes every activity instance, computes the associated intra-case features, and concatenates these features with the tuple representing activity instance $r_{c,i}$ with these features (line 9). [^1] Operator $\oplus$ denotes tuple concatenation.
3.1.3 Inter-Case Features:
Inter-case features capture dependencies across multiple cases, influencing resource contention and case concurrency. Unlike intra-case features, they help predict waiting times and delays. These features are crucial for modeling process dynamics, particularly in resource-constrained systems. Prior research [8] highlights the importance of inter-case features such as work in progress (WIP) and process load (e.g., resource utilization) in capturing dynamic process behavior. Alg. 1 explains the extraction of inter-case features for a given $\mathcal{P}$. It computes three inter-case features: WIP, resource utilization, and the arrival rate of new cases (lines 10-29).
According to Little’s Law [5], the average cycle time ($T_{c}$) of a process is determined by the Work in Progress (WIP) and arrival rate ($\lambda$) via $WIP=\lambda\times T_{c}$. In human-centric processes, cycle time is often dominated by waiting time. We, therefore, hypothesize that features correlated with cycle time can improve waiting time prediction. To compute WIP at time $T^{\text{start}}_{c,k}$, Alg. 1 counts active activity instances in $\mathcal{L}{}$ (lines 11-17). It also estimates the arrival rate as $\lambda=\frac{\text{Total cases}}{\text{Time window}}$ (lines 22-28), using a recent time window (default: 20% of log duration). The set of recent cases is updated dynamically (line 27). Higher WIP and $\lambda$ indicate system congestion, leading to longer expected waiting times.
Queuing theory suggests that waiting times are influenced by resource utilization, the ratio of resource demand to availability [2]. When demand exceeds capacity, utilization reaches 100%, causing delays. Lower demand reduces utilization. Alg. 1 computes utilization by counting active instances per activity at a given time (lines 16-20), helping predictive models account for resource load and capacity. For each trace prefix $\text{$\mathcal{P}$}\subseteq\sigma_{c}$, Alg. 1 computes inter-case features and augments $r_{c,k}$ accordingly.
3.1.4 Sequence Encoding and Construction:
Our approach uses two feature types: intra-case features from individual cases and inter-case features from all active cases. These are encoded into feature vectors for each $\mathcal{P}$, combining both contexts to form input sequences.
Feature Encoding:
To represent the input data for the LSTM, we encode both intra-case and inter-case features in a structured and consistent manner. Intra-case features include attributes specific to each event in a case, such as the activity label, inter-start time, and processing time. Activity labels, being categorical, are embedded using trainable embedding layers rather than one-hot encoding. These embeddings are learned jointly with the LSTM network during supervised training, allowing the model to capture task-specific semantic relationships between activities and efficiently handle a larger set of categories, following the approach in [4]. The temporal features, inter-start time and processing time are treated as continuous variables and normalized to the [0,1] range using min-max normalization based on the training data, as illustrated in Table 2.
Table 2: N-grams snippet for Table. 1.
[…]
1 [0000] 0 5
[0001] 10 5
[0012] 10 40
[0123] 45 5
2 [0000] 0 5
[0001] 10 $\emptyset$
3 [0000] 0 5
[0001] 10 5
[0012] 10 40
Inter-case features, such as WIP, resource utilization, and demand rate ($\lambda$), capture the broader system context. These features are updated at each time step and normalized using min-max scaling for stability. Each feature is aligned with its corresponding event to maintain temporal coherence and provide accurate context at every step.
Sequence Construction:
To construct input sequences and target labels for training, we adopted a fixed-size n-gram approach for each $\mathcal{P}$ in $\mathcal{L}$, as outlined in previous studies [4, 19]. Since variable-length prefixes cannot be directly used as inputs, the n-gram method enables standardization of the temporal dimensionality of the input data hence, capturing sub-sequence patterns of activity instances in $\mathcal{P}$. Each n-gram sequence incorporates both intra-case features (e.g., activity, inter-start time, processing time) and inter-case features (e.g., WIP, resource utilization, $\lambda$), ensuring the model captures both case-specific and system-wide process dynamics. Sequences are generated independently for each $\mathcal{P}$, maintaining consistency across all instances, as shown in Table. 2.
3.1.5 Model Architecture Design and Training:
Existing SM architectures predicts all aspects of the next activity instance in a trace prefix $\mathcal{P}$, i.e., the next activity, its inter-start time, and processing time jointly. However, this joint prediction approach lacks flexibility in scenarios where only partial predictions are required. For example, if the activity and its start time are already known, as is the case for ongoing activities, it is often enough to predict only the processing time. In SM-based architectures, this is difficult because all aspects of the next activity instance are predicted together, and the model cannot easily make just one specific prediction. To address this, we propose a BiLSTM-based Multi-model Predictive Learning Architecture (MM) [7], which leverages bidirectional LSTMs to predict the next activity, inter-start time, and processing time for a trace prefix $\mathcal{P}$. BiLSTM analyzes event sequences bidirectionally, capturing activity dependencies, temporal variations, and overlaps.
MM combines categorical and continuous inputs using pre-trained embeddings, concatenated features, and BiLSTM layers to model sequential dependencies. It follows a modular design with three independent BiLSTM predictors: $\alpha$ for next activity, $\beta$ for inter-start time ($\delta^{\text{start}}_{c,k}$), and $\gamma$ for processing time ($\delta^{\text{proc}}_{c,k}$), as shown in Fig. 1. Each model is trained separately on $\mathcal{L}_{train}$ to capture distinct aspects of the process.
To train the next activity predictor $\alpha$, we extract $\mathcal{P}$’s from $\mathcal{L}_{train}$ and construct inter and intra-case feature vectors using Alg. 1. These features are encoded into input sequences with target labels to train $\alpha$. For the inter-start time predictor $\beta$, we use a similar feature set, but the model additionally takes the actual next activity $a_{c,k+1}$ from $\text{$\mathcal{P}$}_{k+1}$ as input to learn accurate temporal dependencies. Fixed-length input sequences and corresponding labels are generated for training. The processing time predictor $\gamma$ also uses inter- and intra-case features, but focuses on estimating the duration of the next activity $a_{c,k+1}$, incorporating both $a_{c,k+1}$ and its inter-start time $\delta^{\text{start}}_{c,k+1}$. This enables $\gamma$ to capture temporal patterns and case-specific variations for accurate processing time prediction.
3.2 Online Phase
The online phase (Fig. 2) uses the predictive models ($\alpha,\beta,\gamma$) trained during the offline phase to predict case suffixes using a sweep-line-based method, wherein the suffixes of all ongoing cases in the process are predicted in lockstep rather than in isolation. Alg. 2 details the online phase, iteratively processing ongoing cases in chronological order of their predicted start timestamps.
Figure 2: Online Phase
3.2.1 Input:
In the online phase, Alg. 2 takes as input a log $\mathcal{L}$, a cutoff time point $t_{\text{cutoff}}$, and three models $\alpha,\beta,\gamma$. The cutoff time $t_{\text{cutoff}}$ marks the beginning of the online phase, from which future activity instances are predicted in lockstep.
To reflect the system’s state at $t_{\text{cutoff}}$, the log $\mathcal{L}$ is modified by removing all activity instances that start after the $t_{\text{cutoff}}$. For activity instances that begin before but end after $t_{\text{cutoff}}$, the end timestamp is set to $\emptyset$ to indicate the activity was still ongoing. The resulting set of activity instance traces $\sigma_{c}{}$ forms a collection of incomplete trace prefixes, denoted $\text{$\mathcal{P}$}{}_{\emptyset}$, each representing a partial sequence of completed and ongoing activities.
Each $\text{$\mathcal{P}$}{}_{\emptyset}$ is linked to a case identifier and contains the activity name, start timestamp, and if available end timestamp for each instance in $\sigma_{c}{}$, as illustrated in Table 1. These prefixes collectively form the test log ($L^{\prime}$) (line 1). Providing $\text{$\mathcal{P}$}{}_{\emptyset}$ as input to MM offers the historical context required to predict the next activity and its expected start and end timestamps for each ongoing case.
Input: Models $\alpha$, $\beta$, $\gamma$; activity instance log $L$; cutoff $t_{\text{cutoff}}$
Output: Completed log $L^{\prime}$ with predicted suffixes
1
2Let $L^{\prime}\leftarrow\texttt{Extract}(L,t_{\text{cutoff}})$;
3
4Let $t_{\text{sweep}}\leftarrow t_{\text{cutoff}}$;
5
6while $t_{\text{sweep}}\leq\max(T^{\text{start}}\text{ in }L^{\prime})$ do
7  Let $P_{\emptyset}\leftarrow$ $\texttt{Prefix}(L^{\prime},t_{\text{sweep}})$;
8
9 foreach $P_{c}\in P_{\emptyset}$ do
10     foreach $r=(c,a,T^{\text{start}},T^{\text{end}})\in P_{c}$ where $T^{\text{end}}=\emptyset$ do
11        Enhance $P_{c}$ with intra/inter-case features using Alg. 1;
12        Let $\delta^{\text{proc}}\leftarrow\gamma(P_{c})$;
13        $T^{\text{end}}\leftarrow T^{\text{start}}+\delta^{\text{proc}}$;
14        $r\leftarrow(c,a,\delta^{\text{proc}},T^{\text{start}},T^{\text{end}})$;
15
16
17
18 foreach $P_{c}\in P_{\emptyset}$ do
19     Enhance $P_{c}$ with intra/inter-case features using Alg. 1;
20
21    Let $a_{k+1}\leftarrow\alpha(P_{c})$;
22     Let $\delta^{\text{start}}_{k+1}\leftarrow\beta(P_{c},a_{k+1})$;
23     Let $\delta^{\text{proc}}_{k+1}\leftarrow\gamma(P_{c},a_{k+1},\delta^{\text{start}}_{k+1})$;
24
25    $T^{\text{start}}_{k+1}\leftarrow T^{\text{start}}_{k}+\delta^{\text{start}}_{k+1}$;
26     $T^{\text{end}}_{k+1}\leftarrow T^{\text{start}}_{k+1}+\delta^{\text{proc}}_{k+1}$;
27
28    Append $(c,a_{k+1},\delta^{\text{start}}_{k+1},\delta^{\text{proc}}_{k+1},T^{\text{start}}_{k+1},T^{\text{end}}_{k+1})$ to $P_{c}$;
29
30
31 Update $L^{\prime}$ with modified prefixes $P_{c}$;
32
33 $t\leftarrow\left\{T^{\text{start}}\in L^{\prime}\;\middle|\;T^{\text{start}}>t_{\text{sweep}}\text{ and activity}\neq\text{EOT}\right\}$;
34
35 if $t\neq\emptyset$ then
36     $t_{\text{sweep}}\leftarrow\min(t)$
37  else
38     break
return $L^{\prime}$
Algorithm 2 $M$ based Sweep Line Algorithm
To predict suffixes of all ongoing cases in lock step as of $t_{\text{cutoff}}$, the Alg. 2 begins by initializing the $t_{sweep}$ at $t_{\text{cutoff}}$ (line 2), which serves as the starting point for the sweep-line algorithm to begin completing the incomplete trace prefixes ($\text{$\mathcal{P}$}{}_{\emptyset}$). MM integrates three models $\alpha,\beta,\gamma$ in to a sweep-line based simulation algorithm to sequentially predict the next activity, its inter-start time and its processing time for a given incomplete trace prefix $\text{$\mathcal{P}$}{}_{\emptyset}$.
While $t_{sweep}$ is less than the maximum start timestamp in $\mathcal{L}^{\prime}$, the algorithm extracts incomplete trace prefixes ($\text{$\mathcal{P}$}_{\emptyset}$) as of $t_{sweep}$ (line 4). Each $\text{$\mathcal{P}$}_{\emptyset}$ includes traces where the last activity has started but not yet finished. Lines (5-10) handle such cases by first completing these ongoing activities. For each prefix $\text{$\mathcal{P}$}_{c,k-1}$, inter- and intra-case features are extracted using Alg. 1 (line 7) and passed to the prediction model $\gamma$ to estimate the remaining processing time (line 8). The end timestamp is then computed by adding the predicted $\delta^{\text{proc}}_{c,k}{}$ to the start time of $\text{a}_{c,k}$ (line 9). To support concurrent activities, the algorithm applies this prediction to all activity instances in the prefix $P_{c}$ with missing end timestamps, rather than only the most recent one, ensuring consistent suffix prediction in the presence of parallelism.
With complete activity instances, $\text{$\mathcal{P}$}_{c}$ is enhanced with inter & intra-case features using Alg. 1 (line 12). For each enhanced $\text{$\mathcal{P}$}_{c}$ (lines 13-15), MM sequentially applies its models, where $\alpha$ predicts the next activity instance $\text{act}_{c,k+1}$ given trace $\text{$\mathcal{P}$}_{c}$. $\beta$ predicts its inter-start time $\delta^{\text{start}}_{c,k+1}$ given $\text{a}_{c,k+1}$ and $\text{$\mathcal{P}$}_{c}$ and $\gamma$ predicts its processing time given $\text{a}_{c,k+1}$, $\delta^{\text{start}}_{c,k+1}$, and $\text{$\mathcal{P}$}_{c}$.
The start timestamp is obtained by adding $\delta^{\text{start}}_{c,k+1}$ to the previous activity’s start timestamp, while the end timestamp is calculated by adding the processing time to the start timestamp (line 16-17). The predicted activity instances are then appended to $\text{$\mathcal{P}$}_{c}$, extending the ongoing cases (line 18). The updated $\text{$\mathcal{P}$}_{c}$ is subsequently added to $\mathcal{L}^{\prime}$ for the next prediction cycle. The algorithm then advances to the next event time by selecting the minimum start timestamp in $\mathcal{L}^{\prime}$ that is greater than the current and not associated with an end-of-trace (EOT) activity (line 20). If no such timestamp exists, the loop terminates. This process repeats until all cases are completed, meaning the next predicted activity for every trace is EOT.
4 Evaluation
Previous methods use Single Model Architectures (SM), where a single model is trained to predict both the next activity and its timestamp. In contrast, as discussed in Sec. 3, our MM approach applies separate models sequentially to predict the next activity, its inter-start time, and its processing time. This modular design offers greater flexibility, especially when the activity and start time are already known at a given $t_{sweep}$, enabling more accurate processing time predictions. However, it remains to be seen whether this modularity affects predictive performance. The evaluation below addresses the question: How does MM compare to SM in predicting the suffix of ongoing cases in a prefix log?
  1. 1.
    EQ1: In terms of control flow prediction? We hypothesize that MM’s flexibility – enabling the integration of the most suitable control flow prediction model for each event log through targeted training – enables it to outperform SM-based baselines. We also expect that the inter-case features (resource utilization, demand, and capacity) will not enhance control-flow accuracy, as they capture workload properties rather than execution order.
  2. 2.
    EQ2: In terms of inter-start time prediction? Since inter-start times (i.e., time from the start of the current activity to the start of the next activity) are influenced by resource utilization, demand, and capacity, we expect that incorporating inter-case features into our sweep-line approach will improve their prediction. Inter-case features provide needed information on workload and availability, which are determinants of waiting time. Furthermore, we hypothesize that better control flow prediction will further enhance inter-start time estimation. Since activities have varying waiting times, a better next-activity predictor improves inter-start time prediction and hence start timestamp estimation.
  3. 3.
    EQ3: In terms of processing time prediction? Unlike inter-start times, processing times are less dependent on workload factors such as resource utilization, demand and capacity. Instead, they are primarily determined by the nature of the activity itself. Consequently, we do not expect inter-case features to improve processing time prediction. However, we anticipate that a more accurate next-activity prediction could improve processing time estimates. Since activities have distinct processing times, accurate next-activity prediction enhances processing time estimation.
Table 3: Event Log Statistics
Log #Cases #Act-Ins #Act #Act/Case CV Len Avg.Dur(days) Max.Dur(days) CV Dur
BPI17W 30,270 240,854 8 7.96 66.18% 12.66 days 286.07 days 706.70%
BPI12W 8,616 59,302 6 6.88 104.11% 8.91 days 85.87 days 369.44%
INS 1,182 23,141 9 19.58 74.27% 70.93 days 599.9 days 459.36%
ACR 954 4,962 16 5.2 32.13% 14.89 days 135.84 days 157.25%
MP 225 4,503 24 20.01 93.78% 20.63 days 87.5 days 130.40%
CVS 10,000 103,906 15 10.39 11.00% 7.58 days 21.0 days 642.64%
CFS 1,000 21,221 29 26.53 54.60% 0.83 days 4.09 days 83.52%
CFM 2,000 44,373 29 26.57 55.41% 0.76 days 5.83 days 83.82%
P2P 608 9,119 21 15 54.43% 21.46 days 108.31 days 78.04%
Real-Life and Synthetic Datasets:
We evaluate our approach using nine event logs [^2] All logs, including a mix of large- and small-scale processes, are available at supplementary material https://shorturl.at/kJA01, selected for their diverse control-flow structures and temporal characteristics. Table 3 reports for each log the number of cases(#Cases), activity instances (#Act-Ins), and unique activities (#Act). (#Act/Case) is the average number of activities per case. CV Len and CV Dur denotes the coefficient of variation in case length and duration respectively [^3] The coefficient of variation (CV) is computed as: $\text{CV}=\frac{\sigma}{\mu}$, where $\sigma$ is the standard deviation and $\mu$ is the mean of case length or duration.. Avg.Dur and Max.Dur shows average and maximum case durations in days.
All logs include start and end timestamps, as required by our approach. The BPI12W and BPI17W logs capture real-life financial processes from a Dutch institution, with BPI17W being a refined version of BPI12W. [^4] https://doi.org/10.4121/uuid:3926db30-f712-4394-aebc-75976070e91f, https://doi.org/10.4121/uuid:5f3067df-f10b-45da-b98b-86ae4c7a310b These logs represent human-performed activities with moderate process complexity. The INS log represents an insurance claims process and features long case durations and high variability. The ACR log, sourced from a Colombian university’s BPM system, models academic credential recognition, with medium trace lengths and a balanced activity structure. The MP (Manufacturing Production) log captures production operations from an ERP system, characterized by long traces and high activity diversity.
We also include three synthetic logs that simulate real-life operational settings. The CVS log models a retail pharmacy process, based on the simulation in Fundamentals of Business Process Management [5], and serves as a large-scale training set. The CFS and CFM logs are anonymized datasets derived from a confidential process, representing small- and medium-scale versions of the same underlying workflow, both featuring high activity density and resource contention. Finally, the P2P (Purchase-to-Pay) log is a synthetic dataset of a procurement process, offering high structural complexity and longer case durations. Table 3 summarizes the statistics of all logs used in the evaluation.
Measures of Goodness:
To assess sequence similarity between predicted and ground-truth suffixes, we use the Damerau-Levenshtein (DL) distance. DL captures common edit operations – insertions, deletions, substitutions, and transpositions – that may occur between predicted and actual activity sequences [19]. We compute the DL distance for each predicted suffix against its corresponding ground-truth suffix and normalize it by the length of the longer sequence to ensure comparability across traces of varying lengths. A lower normalized DL score indicates higher similarity and better control-flow prediction accuracy.
Moreover, we use the Mean Absolute Error (MAE) to quantify the difference between predicted inter-start times and processing times (relative to the start time of the case) and the actual inter-start and processing times. A higher MAE between a set of predicted timestamps and a set of actual timestamps indicates a lower accuracy (thus lower MAE is better).
Baselines:
The experiments compare the performance of our MM architecture against several variants of a Single Model architecture (SM) baselines adapted from Camargo et al. [4]. Specifically, we evaluate four SM variants: Shared Categorical and Full Shared, each tested with and without inter-case features. In the Shared Categorical variant, only the embeddings of categorical inputs (e.g., activity labels) are shared across tasks, while the remaining layers are task-specific. In contrast, the Full Shared variant employs a fully shared architecture, where both categorical and continuous inputs are processed jointly through the same network layers for all tasks. These configurations allow us to isolate the effects of architectural sharing and context-aware features on predictive performance.
To ensure a fair comparison, all models use the same feature encoding and sequence construction methodology described in (Sec. 3.1). Intra-case features are computed individually per case, while inter-case features are derived by sequentially traversing all cases to capture dynamic workload conditions. Activities are encoded as categorical variables using low-dimensional embeddings, and input sequences are constructed via an n-gram strategy, ensuring standardized and consistent feature representation across models.
Experimental Setup:
To simulate real-life scenarios where models are trained on historical data and applied to ongoing cases, we represent each training and test instance as a pair consisting of a prefix and a suffix trace, denoted as $(\sigma\leq k,\sigma>k)$, where the prefix length $k$ is at least 1. To prevent data leakage, we implement a strict temporal split [23] when dividing the event log into training and testing sets. We define the duration of a log as the time elapsed between the earliest start timestamp (lower bound) and the latest recorded end timestamp (upper bound). The cutoff point $t_{\text{cutoff}}$ is determined as the timestamp where 80% of the total process duration has elapsed. Cases that complete within the first 80% of the timeframe are assigned to the training set, while the remaining 20% are allocated to the test set. Additionally, cases that start before the cutoff but remain ongoing, along with their activity instances, are included in the test set. This ensures the test phase reflects real-world scenarios where predictions handle incomplete cases.
In SM-based experiments, activity instances that start and finish before the cutoff are excluded from the test set, as SM predicts the next activity, inter-start time, and processing time together but cannot handle ongoing instances. In contrast, MM includes ongoing instances since it can predict their processing time given its activity and start timestamp.
Experiments were conducted on a desktop with an NVIDIA RTX 3090 GPU, Intel Core i9 CPU, and 64 GB RAM. Training the three MM models ($\alpha$, $\beta$, $\gamma$) takes 2.5 hours for smaller logs (e.g., ACR, CFS) and 5-6 hours for larger ones (e.g., CVS, BPI2017W). During testing, the sweep-line approach predicts suffixes of ongoing cases at runtime. For BPI2017W, the approach generated 15,000 predicted events in 100 seconds at 150 events per second.
Next Activity Prediction in the MM Architecture:
A core component of the MM architecture is a model $\alpha$ that predicts the next activity in a case. Using an LSTM-based approach, the model generates multiple possible next activities, each assigned a probability score. A sampling method then selects the next activity. We evaluated MM’s accuracy with three sampling methods: Argmax, which selects the most probable activity [19]; Random Choice, which randomly selects an activity based on the probability distribution [4]; and Daemon Action, a heuristic-based method [1]. All three methods performed similarly, but Daemon Action produced the most consistent results across datasets. Therefore, we selected Daemon Action as the preferred sampling method for model $\alpha$.
Table 4: Hyperparameter Optimization
Parameter - LSTM Explanation Search Space
Batch size # of samples to be propagated [32, 64, 128]
Normalization method Preprocessing - Scaling [lognorm, max]
Epochs # of training epochs 200
N_size Size of the n-gram [5, 10, 15, 20, 25]
L_size LSTM layer sizes [50, 100, 150]
Activation Activation function (hidden layers) [selu, tanh]
Optimizer Weight optimizer [Nadam, Adam, SGD, Adagrad]
Hyperparameter Optimization:
We applied hyperparameter optimization within the sets of values in Table 4 for selecting the most suitable three MM models ($\alpha$ for next activity, $\beta$ for inter-start time, $\gamma$ for processing time). Batch size is optimized for computational efficiency and gradient stability, while normalization methods like log normalization and max scaling improve convergence. The n-gram size (N_size) ensures meaningful historical dependencies, and different LSTM layer sizes (L_size) balance model complexity and training efficiency. Activation functions (Selu, Tanh) help maintain stable gradients, and optimizers (Nadam, Adam, SGD, Adagrad) are chosen for optimal convergence. The training set is split (80% training, 20% validation), and MM is trained with diverse hyperparameter combinations. We optimized hyperparameters using a random search over 50 iterations, which provides a practical balance between search space coverage and computational efficiency. Each model was trained for up to 200 epochs with early stopping(patience = 10) to prevent overfitting and reduce unnecessary training time. These settings are consistent with prior studies [4, 1] and were empirically validated to ensure stable convergence. During optimization, we evaluated performance using two metrics: Mean Absolute Error (MAE)  [^5] MAE is defined as $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ for $\delta^{\text{start}}_{c,k}$ and $\delta^{\text{proc}}_{c,k}$ predictions; and Categorical Cross-Entropy Loss [^6] Categorical Cross-Entropy Loss is defined as $\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{n}\sum_{j=1}^{C}y_{ij}\log(\hat{y}_{ij})$ for next-activity prediction.
To evaluate BiLSTM configuration impact, we tested multiple MM variants with varying BiLSTM layer sizes and n-gram window sizes (see Table 4). Larger models marginally improved accuracy but increased training time and reduced throughput. The default configuration (100-unit layer size, 10 n-gram size, 64 batch size, 200 epochs, Adam optimizer) strikes a practical balance between accuracy and efficiency, offering an effective trade-off for predicting case suffixes in business processes.
Results and Interpretation:
We evaluate the proposed architectures using three metrics: DL distance, inter-start time MAE, and processing time MAE. Summary results are shown in the box plots (Fig. 3a, Fig. 4a, and Fig. 5a), with each point representing a log. Detailed per-log results are provided in the corresponding heatmaps (Fig. 3b, Fig. 4b, and Fig. 5b), where bold-bordered cells mark the best-performing architecture per log. Lower values indicate better performance.
Regarding EQ1, we observe in Fig. 3 that MM outperforms SM-based baselines on the control-flow metric. In the heatmap, we see that the outperformance of MM is most visible for event logs with fewer distinct activities (BPI2017W, BPI2012W, CVS, INS, and ACR, which have less than 20 activities). When the number of activities is lower, the embeddings capture richer information about the sequential relations between activity pairs. The next-activity sampling method used in MM is then able to better exploit this information relative to the argmax next-activity sampling method used in the baselines.
(a) Distribution of Damerau-Levenshtein (DL) distance between $M$ and SM.
(b) Heatmap showing DL distances between $M$ and SM across architectures.
Figure 3: DL distance (activity) comparison between Multi-Model ($M$) and Single-Model (SM): (a) box plot, (b) heatmap. Results for EQ1.
(a) Box plot showing MAE of inter-start times for $M$ vs. SM.
(b) Heatmap of MAE in inter-start times between $M$ and SM architectures.
Figure 4: MAE in inter-start time predictions: Multi-Model ($M$) vs. Single-Model (SM): (a) box plot, (b) heatmap. Results for EQ2.
Regarding EQ2, Fig. 4 box plot shows that the three techniques using inter-case features achieve slightly lower inter-case time MAE than the two baselines that rely only on intra-case features. The heatmap in Fig. 4b confirms this observation, particularly for event logs with fewer unique activities (BPI2017W, CVS, INS, and ACR). [^7] An exception is BPI2012W, where all the techniques exhibit similar performance. For logs with larger number of distinct activities, we cannot make a clear conclusion. In the P2P log, the two baselines without inter-case features perform poorly, but in the CFS log, the use of inter-case features does not lead to lower inter-start time MAE and, in fact, MM has the worst performance among all techniques.
Regarding EQ3, Fig. 5 shows that inter-case features have minimal impact on processing time prediction. This is expected since processing time is determined by the activity’s complexity rather than workload, and the inter-case features capture workload information. The heatmap in Fig. 5b confirms this observation. It shows that MM outperforms the baselines in four of the five logs that have less than 20 distinct activities (BPI2012W, CVS, INS, and ACR), but there is no clear pattern in the remaining logs. The outperformance of MM on these four logs is attributable to the fact that it is able to better predict the next-activity, which in turn results in better predictions of the next activity’s processing time.
(a) Box plot showing MAE of processing times for $M$ vs. SM.
(b) Heatmap of MAE in processing times between $M$ and SM architectures.
Figure 5: MAE in processing time predictions: Multi-Model ($M$) vs. Single-Model (SM): (a) box plot, (b) heatmap. Results for EQ3.
The above findings highlight that the use of embeddings has a positive effect in event logs with fewer distinct activities (BPI2017W, BPI2012W, CVS, ACR, INS) but not in logs with 20+ activities. The embeddings we employ (taken from the method in [4]) are based on 3-grams, i.e. they look at which activities occur immediately before or after a given activity. To achieve better accuracy in logs with more distinct activities, we hypothesize that other types of embeddings are needed, e.g. hierarchical classification [9] or adaptive embeddings [18].
Additionally, logs like INS and MP contain longer and more variable traces (mean case length $\approx$ 20 activities, CV of case length >> 65%, with INS having durations from 1 to nearly 600 days and a duration CV of 459%), favoring MM’s modular design in learning decoupled control-flow and timing patterns. MM also excels on BPI2017W despite a shorter mean case length ($\approx$ 7 activities), due to its extremely high duration variability (CV of duration 706%). In contrast, CFS and CFM exhibit short (CV of case length << 60%) and homogeneous traces (mean duration << 1 day, CV of duration 83%), where low variability and limited concurrency make simpler SM architectures competitive.
To quantify the benefit of our proposed approach, we compared MM against the best-performing SM baseline for each metric. MM achieves a 6.89% reduction in DL distance, indicating improved control-flow prediction accuracy. For inter-processing time prediction, MM reduces the mean absolute error (MAE) by 24.44%, and for inter-start time prediction, by 18.35%. These results confirm that MM not only improves predictive performance across all dimensions but does so with substantial error reductions over the most competitive baseline.
5 Conclusion
This paper presents an approach to predict the sequence of remaining activities in ongoing cases of a process, along with the inter-start and processing times. Our proposed approach introduce the concept of predicting suffixes of all ongoing cases in a process, in lockstep, using a sweep line-based approach. This method enables the calculation of inter-case features at each step of the prediction process. These features help in predicting waiting times, as waiting times depend on the overall workload across all ongoing cases. Furthermore, our approach predicts case suffixes using three separate models: one for the next activity, another for inter-start time, and a third for processing time. Our approach enables optimizing each model separately and predicting the end time of ongoing activity instances, given that their activity type and start times.
Our evaluation show that this approach (namely MM) has better accuracy relative to existing baselines that predict the activity and timestamps via one single model (SM approaches). In particular, we observed an improvement in the control-flow metrics, attributable to the fact that we exploited the modularity of the proposed approach to incorporate an optimized method for next-activity sampling. We also observed improvements in inter-start time prediction attributable to the use of inter-case features in the proposed sweep-line method.
In future work, we plan to explore alternative architectures such as Transformers and encoder-decoder models with attention for next activity, inter-start time, and processing time prediction. We also aim to reduce reliance on teacher forcing by mixing predicted and ground-truth activities during training. Additionally, we intend to apply our approach to use cases like capacity planning and scheduling. To further assess the role of inter-case features, we plan to compare our method against state-of-the-art models that do not use such features.
Reproducibility. Source code and Supplementary Material:
https://shorturl.at/RXj4d, https://shorturl.at/kJA01
Acknowledgments. Work funded by the European Research Council (PIX Project) and the Estonian Research Council (PRG1226).